Use the pie chart at the​ right, which shows the number of workers​ (in thousands) by industry for a certain country.

Find the probability that a worker chosen at random was not employed in the mining and construction industry.

A pie chart consists of a circle divided into four sectors with labels as follows: Services, 115,790; Agriculture, forestry, fishing, and hunting, 2931; Manufacturing, 15,728; Mining and construction, 11,102. 115 comma 790 11 comma 102 15 comma 728 2931

A-The probability is . ​(Round to three decimal places as​ needed.)
Answer #1

solution:-
given that total = 115,790+2931+15,728+11,102
= 145551
=> probability that a worker chosen at random was not employed in the mining and construction industry
= 1 - (11,102/145551)
= 1 - 0.076
= 0.924